A deposit of 1 000 $ works with an interest rate of 3%. The deposit's amount, in 2 years time, will be 1 061 $. The total interest earned from this financial placement and the interest earned after one year, are: " with steps"

a.20 $ and 61 $

b. 31 $ and 60 $

c. 61 $ and 30 $

d. 50 $ and 20 $

e. 30 $ and 60 $


1
Expert's answer
2022-02-13T12:12:21-0500

The total interest=1061-1000=61

the interest earned after one year:


"FV1=1000(1+0.03)=1030"

1030-1000=30


c. 61 $ and 30 $